OpenStudy (anonymous):

When you buy 1kg of potatoes at the supermarket, is that 1 kg mass of potatoes OR 1 kg wt of potatoes? Also how many newtons does this weigh?

OpenStudy (anonymous):

1Kg is the mass of potato. weight is analogous of force so its unit is Newton (N). weight=mass*acc. due to gravity=m*g=(1kg)*(9.8m/s^2)=9.8N